The Copper Mining company owns 2 mines, which produce 3 grades of ore - high, medium, and low. The company has a contract to supply a smelting company with 12 tons of high-grade ore, 8 tons of medium-grade ore, and 24 tons of low-grade ore. Each mine produces a certain amount of each type of ore each hour it is in operation. The company has developed the following LP model to determine the number of hours to operate each mine X and Y so that contractual obligations can be met at the lowest cost.




1
Expert's answer
2021-09-16T00:43:56-0400

Solution:

Let x1 = the number of hours that mine 1 operates,

And x2 = the number of hours that mine 2 operates

Constraint functions are:

"x_1 \ge0
\\x_2\ge 0
\\6x_1 + 2x_2 \ge 12\Rightarrow 3x_1+x_2\ge6
\\2x _1+ 2x_2 \ge8\Rightarrow x_1+x_2\ge4
\\4x_1 + 12x_2 \ge 24\Rightarrow x_1+3x_2\ge6"

Minimise cost, "C=200x_1+160x_2"

After introducing slack,surplus,artificial variables:




Similarly, doing this process to 4-iteration,



we get, optimal solution is arrived with value of variables as :

"x_1=1,x_2=3"

Min Z=$680